OpenStudy (anonymous):

Production Planning. A manufacturer builds racing bikes and mountain bikes, with the per unit manufacturing costs shown in the table. The company has budgeted 26,150 for materials and 31,800 for labor. How many bikes of each type can be built? materials labor racing 110 120 mountain 140 180

OpenStudy (anonymous):

Hey :)

OpenStudy (anonymous):

you will make X racing bikes and Y mountain bikes.

OpenStudy (anonymous):

hi :)

OpenStudy (anonymous):

yes, this is where im confused, most of my other problems i pair a number with x and a number with y

OpenStudy (anonymous):

so for each racing bike you will use110 materials and 120 labor, right?

OpenStudy (anonymous):

meaning being, for X number of racing bikes, you will use 110 * X materials and 120 * X labor.

OpenStudy (anonymous):

so 110x + 120y = ??

OpenStudy (anonymous):

140x+180y = ??

OpenStudy (anonymous):

no nono

OpenStudy (anonymous):

it's going to be 110x and 120x

OpenStudy (anonymous):

but you won't get anything from that.

OpenStudy (anonymous):

read this again. "meaning being, for X number of racing bikes, you will use 110 * X materials and 120 * X labor."

OpenStudy (anonymous):

ok so for each racing bike(X) i will charge 230

OpenStudy (anonymous):

haha, noo...

OpenStudy (anonymous):

sorry, i was talking somewhere else. i'm done with that now. let me finish my sentence

OpenStudy (anonymous):

ok for Mountain bikes, Y amount of mountain bikes is going to cost you 140y of materials and 180y of labor.

OpenStudy (anonymous):

now lets look back at the question. we have total of 26150 labor.

OpenStudy (anonymous):

110x + 140y = 26150?

OpenStudy (anonymous):

YES!

OpenStudy (anonymous):

and for the labor?

OpenStudy (anonymous):

120x+180y=31800

OpenStudy (anonymous):

but how do i eliminate those

OpenStudy (anonymous):

now its just algebra

OpenStudy (anonymous):

to make things easier, lets divide each side by 10 for both equations.

OpenStudy (anonymous):

so it makes 11x + 14y = 2615 and 12x + 18y = 3180

OpenStudy (anonymous):

11x+14y=2615 12x+18y=3180

OpenStudy (anonymous):

yes. now we just solve that. it's gonna be a lil messy, i guess.

OpenStudy (anonymous):

do you need help with solving the equations?

OpenStudy (anonymous):

should be 85 racing and 120 mtn bikes

OpenStudy (anonymous):

well that was fast. yeah that does work.

OpenStudy (anonymous):

good job! any question?

OpenStudy (anonymous):

ya but i dont know how that is

OpenStudy (anonymous):

how do i eliminate

OpenStudy (anonymous):

what do 11 and twelve have in common

OpenStudy (anonymous):

132.

OpenStudy (anonymous):

but you wouldn't go for the 11 and 12. that's too much.

OpenStudy (anonymous):

try 14 and 18.

OpenStudy (anonymous):

9 and 7

OpenStudy (anonymous):

and dont forget 2. 9/7/2

OpenStudy (anonymous):

but you probably used your graphing calculator to solve it. i don't think you'll ever need to solve it only with your hands only. so don't worry too much about it.

OpenStudy (anonymous):

thans!!!

OpenStudy (anonymous):

thanks* :P

OpenStudy (anonymous):

no problem. :)